# SVIP: Towards Verifiable Inference of Open-source Large Language Models

### v1

@@ 4.2 SVIP: A Secret-based Protocol for Verifiable LLM Inference
Ultimately, $\tilde{z}^{*}$ is returned to the user to deceive the verification protocol. As shown in our case study in Appendix D.7, this attack achieved an attack success rate (ASR) of $99.90\%$, indicating that the security of the protocol requires further enforcement.
To strengthen the protocol’s security, we introduce a “secret” mechanism. A complete illustration is provided in Figure 2(b). Particularly, the trusted third party assigns a “secret” $s\in\mathcal{S}$ exclusively to the user, which is never shared with the computing provider. Here, $\mathcal{S}$ represents the secret space. Next, we explain how the secret is incorporated into both the self-labeling function and the proxy task.
@@
Property 1 (Secret Distinguishability)
For the same input $x$, given two different secrets $s^{\prime}\neq s$, the resulting labels should be different with high probability:
@@
Property 1 ensures that a malicious computing provider, without access to the specific $s$, cannot determine or naively guess the true label, thus rendering the direct vector optimization attack discussed earlier ineffective. Meanwhile, the user, with knowledge of $s$, can still compute the correct label.
@@
Once the labeling network is optimized, we also need to include the secret $s$ into the proxy task. Inspired by prefix tuning (Li & Liang, 2021), our design is to embed $s$ as a task token using a secret embedding network (e.g., an MLP), denoted as $t_{\psi}(s):\mathcal{S}\rightarrow\mathbb{R}^{d_{\mathcal{M}}}$, parameterized by $\psi$. Note that this secret embedding network $t_{\psi}(s)$ is only kept to the trusted third party. Then, the trusted third party distributes $t_{\psi}(s)$ to the computing provider, who concatenates $t_{\psi}(s)$ with $h_{\mathcal{M}}(x)$, runs $g_{{\theta}}(\cdot)$, and returns $z(x)=g_{{\theta}}(t_{\psi}(s)\oplus h_{\mathcal{M}}(x))$, where $\oplus$ denotes concatenation.
@@ 4.3 Security Analysis
As previously discussed, the direct vector optimization attack described in Eq. (3) is no longer feasible due to the introduction of the secret mechanism. In this section, we discuss other potential attacks as a comprehensive security analysis towards our secret-based protocol. A more detailed discussion of other possible attacks is provided in Appendix C.
@@
To this end, we define the adapter $a_{\lambda}(\cdot):\mathbb{R}^{d_{\mathcal{M}_{alt}}}\rightarrow\mathbb{R}^{d_{\mathcal{M}_{spec}}}$, parameterized by $\lambda$, which transforms the hidden states of $\mathcal{M}_{alt}$ to approximate those of $\mathcal{M}_{spec}$. The returned vector is then $g_{{\theta}^{*}}(t_{\psi^{*}}(s)\oplus a_{\lambda}(h_{\mathcal{M}_{alt}}(x)))$. The attacker’s objective is to minimize the $L_{2}$ distance between the returned vector generated by $\mathcal{M}_{spec}$ and the vector produced by $\mathcal{M}_{alt}$ with the adapter. This can be expressed as:
@@
Once the inverse model is optimized, the true label $y(x,s)$ again becomes accessible to the malicious computing provider. Consequently, the secret-based protocol effectively collapses to the simple protocol without secret protection, leaving it vulnerable to the direct vector optimization attack.
@@
To defend against the attacks discussed above, we propose an update mechanism for our secret-based protocol: (1) In defense of the adapter attack, once the prompt queries for a given secret reach a pre-defined threshold $M^{*}$, the next secret is activated. Meanwhile, we enforce a limit on how often the next secret can be activated, preventing attackers from acquiring too many secrets within a short period. (2) When a total of $N^{*}$ secrets have been used, the entire protocol should be retrained by the trusted third party [^2] Specifically, this retraining can be performed using a different random seed and training recipe. As shown in Section 5.3, the retraining process is efficient.. In practice, the values of $M^{*}$ and $N^{*}$ can be determined empirically, as discussed in Section 5.4.
@@ 5 Experiments
In this section, we evaluate our proposed protocol SVIP through comprehensive experiments to address the following research questions: (1) How accurate is SVIP in verifying whether the specified model is used? (Section 5.2) (2) What are the computational costs associated with SVIP? (Section 5.3) (3) How robust is SVIP against various adaptive attacks? (Section 5.4)
Figure 3: Empirical distribution of the $L_{2}$ distance between the predicted proxy task output $f_{\phi^{*}}(z(x))$ and the label vector $y_{\gamma^{*}}(x,s)$ on the test dataset of LMSYS-Chat-1M. Each figure corresponds to a different specified model. The distributions compare the $L_{2}$ distances when the specified model is used versus various alternative models. The clear separation between the distributions, marked by the vertical threshold line, ensures the high accuracy of our protocol in distinguishing between correct and incorrect model usage.
5.1 Experimental Setup
Datasets and Models
To simulate realistic LLM usage scenarios, we use the LMSYS-Chat-1M conversational dataset (Zheng et al., 2023a), which consists of one million real-world conversations. We filter the dataset to keep only English conversations and extract the user prompts for each conversation. For the models, we select $5$ widely-used LLMs as the specified models, ranging in size from $13$B to $70$B parameters and spanning multiple model families. As alternative models, we use $6$ smaller LLMs, each with parameters up to $7$B. Refer to Appendix D.1 for further details.
Protocol Training Details
The labeling network $y_{\gamma}(\cdot)$ uses a pretrained sentence transformer (Reimers, 2019) to embed the text input $x$ and an MLP to embed the secret $s$, where $s\in\{0,1\}^{d_{s}}$ and $d_{s}$ is set to $48$. The outputs of both embeddings are concatenated and passed through another MLP to produce a continuous label vector of $128$ dimensions. The proxy task feature extractor $g_{\theta}(\cdot)$ is a $4$-layer transformer, while both the proxy task head $f_{\phi}(\cdot)$ and the task embedding network $t_{\psi}(\cdot)$ are implemented as MLPs. Full training details can be found in Appendix D.2.
5.2 Results of Protocol Accuracy
We evaluate the accuracy of our protocol by examining the empirical estimate of FNR and FPR, as outlined in Eq. (16). To apply the verification function in Eq. (8), we first determine the threshold $\eta$ on a validation dataset during proxy task training. We then evaluate the empirical FNR and FPR on a held-out test dataset with $10,000$ samples. For each test prompt, we pair it with $30$ randomly sampled secrets to ensure a reliable evaluation result. For FPR calculations, we simulate scenarios where the computing provider uses an alternative, smaller LLM to produce the hidden representations, and applies $g_{\theta^{*}}(\cdot)$ on those outputs [^3] If the hidden dimension of the alternative LLM, $d_{\mathcal{M}_{alt}}$, differs from that of the specified model, $d_{\mathcal{M}_{spec}}$, we apply a random projection matrix $W\in\mathbb{R}^{d_{\mathcal{M}_{alt}}\times d_{\mathcal{M}_{spec}}}$ to align the dimensions, where each element of $W$ is sampled from a standard normal distribution.. Additionally, we implement a Random baseline where the computing provider generates random hidden representations directly without using any LLM.
As shown in Table 1, SVIP consistently achieves low FNR and FPR across all specified LLMs, demonstrating its effectiveness in verifying whether the correct model is used. The FNR remains below $5\%$, indicating that our protocol rarely falsely accuses an honest computing provider. Moreover, when faced with a dishonest provider, the FPR stays under $3\%$ regardless of the alternative model employed, highlighting the protocol’s strong performance in detecting fraudulent behavior. Figure 3 shows the empirical test distribution of $d(x,s)$, the $L_{2}$ distance between the predicted proxy task output and the label vector, under different model usage scenarios. The clear separation in the distributions provides strong evidence for the high accuracy of SVIP: when the specified model is actually used, $d(x,s)$ is significantly smaller compared to when an alternative model is used.
Table 1: FNR and FPR across different specified models on the test dataset of LMSYS-Chat-1M. Our protocol keeps FNR under $5\%$ and FPR under $3\%$ across all scenarios.
@@ 5.2 Results of Protocol Accuracy
Evaluation on Unseen Dataset
To assess the generalizability of our protocol, we evaluate its accuracy on unseen datasets using the proxy task model and threshold initially trained on the LMSYS-Chat-1M dataset. Specifically, we test on the ToxicChat dataset (Lin et al., 2023), which contains toxic user prompts that were not included in the training stage, representing a reasonable level of distribution shift. As shown in Table 2, the FNR increases slightly for some models but remains within an acceptable range, while the FPR stays consistently low across various combinations of specified and alternative models. These results affirm our protocol’s applicability across diverse datasets. Additional results on the web_questions dataset (Kwiatkowski et al., 2019), which contains popular questions from real users, are provided in Appendix D.3.
Table 2: FNR and FPR on the ToxicChat dataset. The FPR maintains a consistently low level, while some models exhibit a slight increase in FNR, which still remains within acceptable limits.
Specified Model FNR $\downarrow$ FPR $\downarrow$
Llama-2-13B 3.40% 4.33% 3.65% 3.24% 4.21% 4.53% 5.12% 4.50%
GPT-NeoX-20B 15.35% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
OPT-30B 2.56% 0.00% 0.08% 0.12% 0.06% 0.18% 0.02% 0.04%
Falcon-40B 10.30% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Llama-3.1-70B 9.24% 4.40% 5.83% 5.51% 6.12% 6.47% 5.27% 5.36%
5.3 Computational Cost Analysis of the Protocol
During the deployment stage, a practical protocol should introduce minimal computational cost for both the computing provider and the user, specifically in terms of runtime and GPU memory usage. Table 3(a) details the runtime per query and GPU memory consumption. Across all specified models, the verification process takes under $0.01$ seconds per query for both the computing provider and the user. For example, verifying the Llama-2-13B model for each query takes only $0.0017$ seconds for the computing provider and $0.0056$ seconds for the user, in stark contrast to zkLLM (Sun et al., 2024), where generating a single proof requires $803$ seconds and verifying the proof takes $3.95$ seconds for the same LLM. The proxy task feature extractor $g_{\theta}(\cdot)$, run by the computing provider, consumes approximately $980$ MB of GPU memory, imposing only minimal overhead. On the user side, the proxy task head $f_{\phi}(\cdot)$ and labeling network $y_{\gamma}(\cdot)$ require a total of $1428$ MB, making it feasible for users to run on local machines without high-end GPUs. Additionally, we record the required proxy task retraining time in Table 3(b). Overall, retraining the proxy task takes less than $1.5$ hours on a single GPU, allowing for efficient protocol update.
Table 3: Computational costs of SVIP. All measurements were recorded on a single NVIDIA L40S GPU. (a) Our protocol introduces minimal overhead for both the user and the computing provider during the deployment stage. (b) Retraining the proxy task is computationally affordable.
(a) Deployment stage costs.
Specified Model Runtime (Per Query) GPU Memory Usage
User Computing Provider User Computing Provider
Llama-2-13B 0.0056 s 0.0017 s
GPT-NeoX-20B 0.0057 s 0.0017 s
OPT-30B 0.0057 s 0.0018 s 1428 MB 980 MB
Falcon-40B 0.0057 s 0.0018 s
Llama-3.1-70B 0.0057 s 0.0019 s
(b) Proxy task retraining costs.
Specified Model Proxy Task Retraining Time
Llama-2-13B 4492 s
GPT-NeoX-20B 4500 s
OPT-30B 4580 s
Falcon-40B 4596 s
Llama-3.1-70B 5125 s
Figure 4: Attack Success Rate for the adapter attack on the test dataset of LMSYS-Chat-1M, plotted as a function of the number of prompt samples collected under each single secret.
5.4 Results of Protocol Security
Robustness Evaluation Against Adapter Attack
To simulate the adapter attack, we assume an attacker collects a dataset of size $M$, consisting of prompt samples associated with a single secret $s$. The attack follows the optimization process outlined in Eq. (9) , and is considered successful if the resulting adapter passes the verification function when secret $s$ is applied [^4] Specifically, the attack succeeds if: $\|f_{\phi^{*}}(g_{{\theta}^{*}}(t_{\psi^{*}}(s)\oplus a_{{\lambda}^{*}}(h_{\mathcal{M}_{alt}}(x)))-y_{\gamma^{*}}(x,s)\|_{2}\leq\eta$.. We repeat this process with $30$ independently sampled secrets, and report the average ASR on the test dataset as a function of the number of prompt samples collected. The experiment is conducted with $3$ specified LLMs, each paired with $3$ smaller alternative models. Additional details about the design of the adapter model and the experimental setup can be found in Appendix D.5.
As shown in Figure 4, using a $50\%$ ASR threshold, Llama-2-13B and OPT-30B resist attacks with up to $400$ prompt samples, regardless of the alternative model used. For Llama-3.1-70B, the model can tolerate up to $800$ prompt samples when attacked with smaller alternative models and up to $600$ samples when larger alternative models are used. Based on these results, we recommend setting $M^{*}$, the maximum number of prompt queries allowed under a single secret before a new secret is activated, in the range from $400$ to $600$, depending on the specified LLM.
@@ 5.4 Results of Protocol Security
In this attack scenario, we assume the attacker has collected $N$ secret-embedding pairs and uses a $3$-layer MLP as the inverse model to predict the original secret from its embedding. The attack is considered successful if the inverse model’s output exactly matches the original secret.
Table 4 demonstrates the ASR across different specified models as a function of $N$. The attacker is unable to recover any secrets when $N\leq 10,000$. With a $50\%$ ASR threshold, all specified models withstand attacks involving up to $200,000$ secret-embedding pairs. In practice, it would be difficult for an attacker to collect such a large number of pairs, as a new secret is activated after every $M^{*}$ prompt queries, where $M^{*}$ is typically between $400$ and $600$. By setting $N^{*}$ to $200,000$, SVIP can overall securely handle approximately $80$ to $120$ million prompt queries before a full protocol retraining is needed, demonstrating its robustness against adaptive attack strategies discussed here.
@@ 6 Conclusion
In this paper, we formalize the problem of verifiable inference in the context of LLMs. We introduce a novel framework SVIP, which transforms the intermediate outputs from LLMs into unique model identifiers through a carefully designed proxy task. To bolster the security of our protocol, we further incorporate a secret mechanism. We also provide a thorough analysis of potential attack scenarios. Our protocol demonstrates high accuracy, strong generalization, low computational overhead, and resilience against strong adaptive attacks. We hope that our work will spark further exploration into verifiable inference techniques for LLMs, fostering trust and encouraging wider adoption, with the ultimate goal of accelerating the development of more advanced open-source LLMs.

### v2

@@ 3.2 SVIP: A Secret-based Protocol for Verifiable LLM Inference
Ultimately, $\tilde{z}^{*}$ is returned to the user to deceive the verification protocol. As shown in Appendix F.8, this attack achieved an attack success rate (ASR) of $99.90\%$, indicating that the protocol’s security requires further enforcement.
To strengthen the protocol’s security, we introduce a “secret" mechanism. A complete illustration is provided in Figure 3(b). Particularly, the platform assigns a “secret” $s\in\mathcal{S}$ exclusively to the user, which is never shared with the computing provider. Here, $\mathcal{S}$ represents the secret space. For example, $\mathcal{S}$ can be defined as the space of $d_{s}$-dimensional binary vectors, represented as $\{0,1\}^{d_{s}}$.
@@
Property 1 (Secret Distinguishability).
For the same input $x$, given two different secrets $s^{\prime}\neq s$, with a pre-defined lower-bound probability $\delta$, the resulting labels should be different with high probability:
@@
Property 1 ensures that a malicious computing provider, without access to the specific $s$, cannot determine or naively guess the true label, thus rendering the direct vector optimization attack ineffective. Meanwhile, the user, with knowledge of $s$, can still compute the correct label.
@@
Once the labeling network is optimized, we also need to include the secret $s$ into the proxy task. Our design is to embed $s$ as a task token using a secret embedding network (e.g., an MLP), denoted as $t_{\psi}(s):\mathcal{S}\rightarrow\mathbb{R}^{d_{\mathcal{M}}}$, parameterized by $\psi$. Note that this secret embedding network $t_{\psi}(s)$ is only kept to the platform. Then, the platform distributes $t_{\psi}(s)$ to the computing provider, who concatenates $t_{\psi}(s)$ with $h_{\mathcal{M}}(x)$, runs $g_{{\theta}}(\cdot)$, and returns $z(x)=g_{{\theta}}(t_{\psi}(s)\oplus h_{\mathcal{M}}(x))$, where $\oplus$ denotes concatenation.
@@ 3.3 Security Analysis
As previously discussed, the direct vector optimization attack described in Eq. (3) is no longer feasible due to the introduction of the secret mechanism. In this section, we discuss other potential attacks as a security analysis towards our protocol. Additional possible attacks are discussed in Appendix E.
@@
To this end, we define the adapter $a_{\lambda}(\cdot):\mathbb{R}^{d_{\mathcal{M}_{alt}}}\rightarrow\mathbb{R}^{d_{\mathcal{M}_{spec}}}$, parameterized by $\lambda$, which transforms the hidden states of $\mathcal{M}_{alt}$ to approximate those of $\mathcal{M}_{spec}$. The returned vector is then $g_{{\theta}^{*}}(t_{\psi^{*}}(s)\oplus a_{\lambda}(h_{\mathcal{M}_{alt}}(x)))$. The attacker’s objective is to minimize the $L_{2}$ distance between the returned vector generated by $\mathcal{M}_{spec}$ and the vector produced by $\mathcal{M}_{alt}$ with the adapter:
@@
Once the inverse model is optimized, the true label $y(x,s)$ again becomes accessible to the malicious provider. Consequently, the secret-based protocol effectively collapses to the simple protocol without secret protection, leaving it vulnerable to the direct vector optimization attack.
@@
To defend against the attacks discussed above, we propose an update mechanism for our secret-based protocol: (1) In defense of the adapter attack, once the prompt queries for a given secret reach a pre-defined threshold $M^{*}$, the next secret is activated. Meanwhile, we enforce a limit on how often the next secret can be activated, preventing attackers from acquiring too many secrets within a short period. (2) When a total of $N^{*}$ secrets have been used, the entire protocol should be retrained by the platform [^3] Specifically, this retraining can be performed using a different random seed and training recipe. As shown in Section 4.2, the retraining process is efficient.. In practice, the values of $M^{*}$ and $N^{*}$ can be determined empirically, as discussed in Section 4.3.
@@ 4 Experiments
Experiment Setup
To simulate realistic LLM usage scenarios, we primarily use the LMSYS-Chat-1M conversational dataset [55], which consists of one million real-world conversations. Results on additional datasets are provided in Appendix F.3. For the models, we select $5$ widely-used LLMs as the specified models, ranging in size from $13$B to $70$B parameters and spanning multiple model families. As alternative models, we use $6$ smaller LLMs, each with parameters up to $7$B. Refer to Appendix F.1 for details. The labeling network $y_{\gamma}(\cdot)$ uses a pretrained sentence transformer [37] to embed the text input $x$ and an MLP to embed the secret $s$, where $s\in\{0,1\}^{d_{s}}$ and $d_{s}$ is set to $48$. The outputs of both embeddings are concatenated and passed through another MLP to produce a continuous label vector of $128$ dimensions. The proxy task feature extractor $g_{\theta}(\cdot)$ is a $4$-layer transformer, while both the proxy task head $f_{\phi}(\cdot)$ and task embedding network $t_{\psi}(\cdot)$ are implemented as MLPs. Full details can be found in Appendix F.2.
Table 2: FNR and FPR on the test dataset of LMSYS-Chat-1M. Our protocol keeps FNR under $5\%$ and FPR under $3\%$ across all scenarios. We implement a Random baseline where the computing provider generates random hidden representations directly without using any LLM.
@@
4.1 Results of Protocol Accuracy
We evaluate the accuracy of our protocol by examining the empirical estimate of FNR and FPR, as outlined in Eq. (14). To apply the verification function in Eq. (8), we first determine the threshold $\eta$ on a validation dataset during proxy task training. We then evaluate the empirical FNR and FPR on a held-out test dataset with $10,000$ samples. For each test prompt, we pair it with $30$ randomly sampled secrets to ensure a reliable evaluation result. For FPR calculations, we simulate scenarios where the computing provider uses an alternative, smaller LLM to produce the hidden representations, and applies $g_{\theta^{*}}(\cdot)$ on those outputs.
As shown in Table 2, SVIP consistently achieves low FNR and FPR for individual queries across all specified LLMs. The FNR remains below $5\%$ per query, indicating that our protocol rarely falsely accuses an honest computing provider. Moreover, when faced with a dishonest provider, the FPR stays under $3\%$ per query regardless of the alternative model employed, highlighting the protocol’s strong performance in detecting fraudulent behavior. We further demonstrate the generalizability of our protocol to unseen datasets in Appendix F.3.
Figure 4 shows the empirical test distribution of $d(x,s)$, the $L_{2}$ distance between the predicted proxy task output and the label vector, under different model usage scenarios. The clear separation in the distributions provides strong evidence for the high accuracy of SVIP: when the specified model is actually used, $d(x,s)$ is significantly smaller compared to when an alternative model is used.
Figure 4: Empirical distribution of the $L_{2}$ distance between the predicted proxy task output $f_{\phi^{*}}(z(x))$ and the label vector $y_{\gamma^{*}}(x,s)$ on the test dataset of LMSYS-Chat-1M. Each figure corresponds to a different specified model. The distributions compare the $L_{2}$ distances when the specified model is used versus various alternative models. The clear separation between the distributions, marked by the vertical threshold line, ensures the high accuracy of our protocol in distinguishing between correct and incorrect model usage. More examples can be found in Figure 8.
@@ 4.1 Results of Protocol Accuracy
A Hypothesis Testing Framework for a Single Final Conclusion
In practical scenarios, conclusions about a computing provider’s honesty are based on multiple different prompt queries rather than a single one. A hypothesis testing framework can be adopted to combine the results of each individual query and reach a single final conclusion. With FPRs and FNRs below $5\%$ for each individual query, the user can draw a final conclusion about the provider’s honesty with high confidence. For instance, by employing only $30$ different queries, the type-I and type-II error rates of the final conclusion are effectively driven to near zero, demonstrating the strong robustness of SVIP.
As an illustrative case, when using Llama-3.1-70B as the specified model and Llama-2-7B as the alternative, we achieve an FPR of $0.81\%$ and an FNR of $3.13\%$. With a properly chosen decision threshold, the type-I error rate (incorrectly flagging an honest provider as dishonest) and type-II error rate (failing to detect a dishonest provider) rates are $1.7\times 10^{-49}$ and $0.0$, respectively. Refer to Appendix C.2 for detailed analysis and results.
4.2 Computational Cost Analysis of the Protocol
Table 12(a) details the runtime per prompt query and GPU memory consumption during the deployment stage. Across all specified models, the verification process takes under $0.01$ seconds per prompt query for both the computing provider and the user. For example, verifying the Llama-2-13B model for each prompt query takes only $0.0017$ seconds for the computing provider and $0.0056$ seconds for the user, in stark contrast to zkLLM [40], where generating a single proof requires $803$ seconds and verifying the proof takes $3.95$ seconds for the same LLM. The proxy task feature extractor $g_{\theta}(\cdot)$, run by the computing provider, consumes approximately $980$ MB of GPU memory, imposing only minimal overhead. On the user side, the proxy task head $f_{\phi}(\cdot)$ and labeling network $y_{\gamma}(\cdot)$ require a total of $1428$ MB, making it feasible for users to run on local machines without high-end GPUs. Additionally, we record the required proxy task retraining time in Table 12(b). Overall, retraining the proxy task takes less than $1.5$ hours on a single GPU, allowing for efficient protocol update.
4.3 Results of Protocol Security
Robustness Evaluation Against Adapter Attack
To simulate the adapter attack, we assume an attacker collects a dataset of size $M$, consisting of prompt samples associated with a single secret $s$. The attack is considered successful if the resulting adapter passes the verification function when secret $s$ is applied. Additional details about the experimental setup can be found in Appendix F.6.
As shown in Figure 5, using a $50\%$ ASR threshold, Llama-2-13B resist attacks with up to $400$ prompt samples, regardless of the alternative model used. For Llama-3.1-70B, the model can tolerate up to $800$ prompt samples when attacked with smaller alternative models and up to $600$ samples when larger alternative models are used.
Figure 5: Attack Success Rate for the adapter attack, plotted as a function of the number of prompt samples collected under each single secret.
@@ 4.3 Results of Protocol Security
We assume the attacker has collected $N$ secret-embedding pairs to train an inverse model to predict the original secret from its embedding. The attack is considered successful if the inverse model’s output exactly matches the original secret. Table 3 demonstrates the ASR across different specified models as a function of $N$. The attacker is unable to recover any secrets when $N\leq 10,000$. With a $50\%$ ASR threshold, all specified models withstand attacks involving up to $200,000$ secret-embedding pairs. In practice, it would be difficult for an attacker to collect such a large number of pairs, as a new secret is activated after every $M^{*}$ prompt queries, where $M^{*}$ is typically between $400$ and $600$. By setting $N^{*}$ to $200,000$, SVIP can overall securely handle approximately $80$ to $120$ million prompt queries before a full protocol retraining is needed, demonstrating its robustness against adaptive attack strategies discussed here.
@@ 5 Conclusion
In this paper, we present SVIP, a novel framework that enables accurate, efficient, and robust verifiable inference for LLMs. We hope that our work will spark further exploration into this area, fostering trust and encouraging wider adoption of open-source LLMs.